OpenStudy (azureilai):

Can someone tell me when we learn about Parametric functions? like what level of math class?

OpenStudy (anonymous):

I'd say most students are introduced to parametric functions in pre-calc or at the calculus level. In any case, you probably already have been introduced to them, you just haven't realized it. A function like \(y=f(x)\) is a parametric function, the parameter being \(x\). The one's you may be thinking of have the form \(\begin{cases}y=f(t)\\x=g(t)\end{cases}\), with parameter \(t\). The functions you're familiar with would be expressed as \(\begin{cases}y=f(x)\\x=x\end{cases}\).